# for Unseen Language Generalization

### v4

@@ 3 Unseen Languages Adaptation with LinguAlchemy
In this section, we provide an overview of how LinguAlchemy can capture linguistic constraint and how is the intuition behind LinguAlchemy. We also discuss in detail how do we align model representations with the linguistic vector.
@@ 3.1 Does Multilingual LMs capture Linguistic Constraints?
In this work, we define the linguistic knowledge as a vector gathered from URIEL vector (Littell et al., 2017). We chose three distinct linguistic knowledge from the database, namely ’syntax_knn’, ’syntax_average’ [^1] In this work, we chose the ’knn’ and ’average syntax features. These include consensus values (like averages) and predicted values (such as kNN regressions based on phylogenetic or geographical neighbors), and ’geo’ features. The choice of ’syntax_knn’ and ’syntax_average’ is motivated by the typological nature of syntax. Syntax in languages varies widely; hence, by using aggregate measures like averages and k-nearest neighbors (kNN), we can capture a more general representation of syntactic features across languages. Note that in our experiments, we excluded phonological features and language family attributes from our analysis as they are less relevant to textual data and have limited granularity for understanding linguistic variations.
@@ 3.2 Proof of Concept
We investigate whether multilingual language models (MLMs) like mBERT Devlin et al. (2018) can capture linguistic constraints by aligning mBERT language embeddings with URIEL vectors to assess how they represent seen and unseen languages. This includes examining how well mBERT’s embeddings correspond to the typological and geographical features detailed in URIEL. In Figure 2, sentence embeddings (green dots) from mBERT, derived from the last hidden state of multilingual training data, and URIEL vectors (brown dots)—structured representations from the URIEL database—are projected into the same space. A matrix $W$ is used to linearly project sentence embeddings, minimizing the mean squared error with URIEL vectors. This alignment is showcased in Figure 2 using UMAP for dimensionality reduction for visualization purpose.
Figure 2 presents a visual analysis facilitated by UMAP McInnes et al. (2018), showing the correlation between mBERT language representation and the linguistic vectors from the URIEL database ($R^{2}$ = 0.816). By leveraging UMAP, the plot highlight the principal variances within the joint feature space of the embeddings and vectors. The spatial representation of languages on this plot mirrors their linguistic and geographical relatedness, as encapsulated by mBERT. This visualization shows the model’s ability to mirror linguistic typologies, with languages sharing common roots such as ’de-DE’ and ’nl-NL’ naturally clustering together. The density and arrangement of these clusters potentially reflect mBERT capacity to capture and represent language family traits. Conversely, the presence of sparser clusters or outliers requires us to carefully check mBERT’s coverage and consistency in representing different linguistic features. We also formally defined the language representation alignment in the Appendix B.
@@
Table 1 shows the linguistic vector ablation experiment conducted on two models, mBERT and XLM-R. The vector ablation experiment demonstrates that the combination of syntax_knn, syntax_avg, and geographical features yields the best results. Therefore, we will use this combination of features in subsequent experiments.
@@ 4 Experiment Setting
Our research employs two state-of-the-art language models: Multilingual BERT Base (mBERTBase) and XLM-RoBERTa Base (XLM-RBase). In our training process, we use a learning rate of $5\times 10^{-5}$, train for 30 epochs, and measure performance based on accuracy for MASSIVE and MasakhaNews, and Pearson correlation for SemRel. Each training takes at most 5 hours using a single A100 GPU.
@@ 5.1 LinguAlchemy Performance
Despite the variations in performance, the potential of LinguAlchemy is particularly clear in scenarios where zero-shot performance is inherently weak. Our hypothesis is that the model indirectly leverages familiar scripts encountered during pretraining and helps its ability to effectively handle UNK tokens. Advances in models using byte-level tokenization units theoretically reduce or eliminate OOV tokens; however, our evaluations across the MASSIVE, MasakhaNews, and SemRel2024 datasets, as shown in Table 3, confirm that UNK tokens have a minimal impact, demonstrating the robustness of LinguAlchemy in such environments. For contexts where UNK token rates are high, the solution might be orthogonal to our approach, requiring further improvement in the base models or tokenizers that could later be integrated with LinguAlchemy.
@@ 5.2 Effect of Scaling URIEL loss
The classification and URIEL losses are not on the same scale. Therefore, simply adding both losses together means that the model will give more weight to the loss with the higher magnitude. When observing both the classification and URIEL losses during the early stages of training, we note that the classification loss is around 10 times larger than the URIEL loss. In this part, we explore the effect of different scaling factors for the URIEL loss.
Figure 3: Average performance of unseen languages under various URIEL loss scales.
@@
We explore consistently scaling up the URIEL loss across various scaling factors. The results can be seen in Figure 3. It is important to note that, due to the use of the scale-invariant optimizer AdamW, there is no concern about gradients becoming too large as a result of the high losses. Generally, we observe that a scaling factor of 10x slightly outperforms other scaling factors, and the performance appears to decline with higher scale factors.
[…]
@@
One issue with introducing a scaling factor is the addition of another tunable hyperparameter. Intuitively, we aim for a balanced weight between the classification and URIEL losses. Therefore, instead of expensively testing different scaling factors, an adaptive scaling factor is more cost-effective and beneficial. Here, we explore two ideas: dynamic and trainable factors. The results of these approaches can be seen in Table 6.
@@
Interestingly, these dynamic scale factors do not significantly outperform a constant factor. In contrast, a 10x scaling achieves the best performance in mBERT, while dynamic scaling barely outperforms the 10x scaling in XLM-R. Therefore, in a limited budget scenario, a suggested 10x scaling factor should suffice, and one may explore different scaling factors given more computational resources.
@@ 6 Conclusion
We introduced LinguAlchemy, a novel approach that demonstrates strong performance across 27 unseen languages in a 60-class intent classification task. Our method hinges on the integration of linguistic knowledge through the URIEL vectors, enhancing the language model’s ability to generalize across a diverse set of languages. We also proposed AlchemyScale and AlchemyTune, which employs a hyperparameter search for the URIEL scaling factor. This is achieved by two key strategies: (1) weight-averaging classification and URIEL loss, and (2) learning to balance the scale between classification and URIEL loss, thus ensuring a more adaptable and robust performance. LinguAlchemy achieves a massive performance improvement on low-resource languages in multiple downstream tasks including intent classification ($\uparrow 18.50$), news classification ($\uparrow 31.46$), and semantic relatedness ($\uparrow 0.32$).
@@ 4 Sino-Tibetan Train Group 3 + my-MM, zh-CN, zh-TW 36
Appendix B Algorithm
Algorithm 1 Language Representation and Alignment Process
0:  Aligned Language Representations $A$
  $F_{U}\leftarrow$ ExtractFeatures($U$)
  $S\leftarrow\{\}$
  for each sentence $x$ in $D$ do
     $H_{x}\leftarrow$ GetLastHiddenStates($x$, $\Theta$)
     $N_{x}\leftarrow$ CountAttentionMasked($x$)
     $R_{x}\leftarrow\frac{\textsc{Sum}(H_{x})}{N_{x}}$
     $S\leftarrow S\cup\{R_{x}\}$
  end for
  $W,b\leftarrow$ InitializeParameters()
  for each training epoch do
     $P_{U}\leftarrow(W\times S)+b$
     $loss\leftarrow$ ComputeLoss($P_{U}$, $F_{U}$)
     $W,b\leftarrow$ UpdateParametersWithConstraint($W,b,loss$)
  end for
  $A\leftarrow\{\}$
  for each sentence representation $s$ in $S$ do
     $A_{m}\leftarrow(W\times s)+b$
     $A\leftarrow A\cup\{A_{m}\}$
  end for
Appendix C Language Family Experiment
@@
Appendix D Appendix: Language Identification (LID) Experiments

### v6

@@ 3 Unseen Languages Adaptation with LinguAlchemy
In this section, we provide an overview of how LinguAlchemy captures linguistic constraints and explain the intuition behind it. We also discuss in detail how we align model representations with the linguistic vectors.
@@ 3.1 Does Multilingual LMs capture Linguistic Constraints?
We define the linguistic knowledge as a vector gathered from URIEL vector (Littell et al., 2017). We chose three distinct linguistic knowledge from the database, namely ’syntax_knn’, ’syntax_average’ [^1] In this work, we chose the ’knn’ and ’average syntax features. These include consensus values (like averages) and predicted values (such as kNN regressions based on phylogenetic or geographical neighbors), and ’geo’ features. The choice of ’syntax_knn’ and ’syntax_average’ is motivated by the typological nature of syntax. Syntax in languages varies widely; hence, by using aggregate measures like averages and k-nearest neighbors (kNN), we can capture a more general representation of syntactic features across languages. Note that in our experiments, we excluded phonological features and language family attributes from our analysis as they are less relevant to textual data and have limited granularity for understanding linguistic variations.
@@ 3.2 Proof of Concept
We investigate whether PLMs like mBERT Devlin et al. (2018) can capture linguistic constraints by aligning mBERT language embeddings with URIEL vectors to assess how they represent seen and unseen languages. This includes examining how well mBERT’s embeddings correspond to the typological and geographical features detailed in URIEL. In Figure 2, sentence embeddings (green dots) from mBERT, derived from the last hidden state of multilingual training data, and URIEL vectors (brown dots)—structured representations from the URIEL database—are projected into the same space. A matrix $W$ is used to linearly project sentence embeddings, minimizing the mean squared error with URIEL vectors. This alignment is showcased in Figure 2 using UMAP for visualization purpose.
Figure 2 presents a visual analysis facilitated by UMAP McInnes et al. (2018), showing the correlation between mBERT language representation and the linguistic vectors from the URIEL database ($R^{2}$ = 0.816). By leveraging UMAP, the plot highlight the principal variances within the joint feature space of the embeddings and vectors. The spatial representation of languages on this plot mirrors their linguistic and geographical relatedness, as encapsulated by mBERT. This visualization shows the model’s ability to mirror linguistic typologies, with languages sharing common roots such as ’de-DE’ and ’nl-NL’ naturally clustering together. The density and arrangement of these clusters potentially reflect mBERT capacity to capture and represent language family traits. Conversely, the presence of sparser clusters or outliers requires us to carefully check mBERT’s coverage and consistency in representing different linguistic features. We also formally defined the language representation alignment in Algorithm 1.
@@
Table 1 shows the results of the linguistic vector ablation experiment for mBERT and XLM-R, testing different feature combinations. For mBERT, individual features like syntax_avg and syntax_knn perform similarly ( 66.41%), with only minor improvement (66.47%) when combined with geo. In contrast, XLM-R benefits more from feature combinations, achieving its highest accuracy (80.80%) when syntax_knn, syntax_avg, and geo are combined. Even individually, syntax_avg (80.76%) performs well for XLM-R, highlighting the model’s stronger ability to leverage syntactic information. These results suggest that combining syntax and geographical features yields optimal performance, especially for XLM-R, and this combination will be used in subsequent experiments.
@@
Algorithm 1 Language Representation and Alignment Process
0:  Aligned Language Representations $A$
  $F_{U}\leftarrow$ ExtractFeatures($U$)
  $S\leftarrow\{\}$
  for each sentence $x$ in $D$ do
     $H_{x}\leftarrow$ GetLastHiddenStates($x$, $\Theta$)
     $N_{x}\leftarrow$ CountAttentionMasked($x$)
     $R_{x}\leftarrow\frac{\textsc{Sum}(H_{x})}{N_{x}}$
     $S\leftarrow S\cup\{R_{x}\}$
  end for
  $W,b\leftarrow$ InitializeParameters()
  for each training epoch do
     $P_{U}\leftarrow(W\times S)+b$
     $loss\leftarrow$ ComputeLoss($P_{U}$, $F_{U}$)
     $W,b\leftarrow$ UpdateParametersWithConstraint($W,b,loss$)
  end for
  $A\leftarrow\{\}$
  for each sentence representation $s$ in $S$ do
     $A_{m}\leftarrow(W\times s)+b$
     $A\leftarrow A\cup\{A_{m}\}$
  end for
@@ 4 Experiment Setting
Our study employs two widely used and resource-efficient multilingual language models: Multilingual BERT Base (mBERTBase) and XLM-RoBERTa Base (XLM-RBase). In our training process, we use a learning rate of $5\times 10^{-5}$, train for 30 epochs, and measure performance based on accuracy for MASSIVE and MasakhaNews, and Pearson correlation for SemRel. Each training takes at most 5 hours using a single A100 GPU.
@@ 5.1 LinguAlchemy Performance
Despite the variations in performance, the potential of LinguAlchemy is particularly clear in scenarios where zero-shot performance is inherently weak. Our hypothesis is that the model indirectly leverages familiar scripts encountered during pretraining and helps its ability to effectively handle UNK tokens. Advances in models using byte-level tokenization units theoretically reduce or eliminate OOV tokens; however, our evaluations across the MASSIVE, MasakhaNews, and SemRel2024 datasets, as shown in Table 3, confirm that UNK tokens have a minimal impact, thus showing the robustness of LinguAlchemy in such environments. For contexts where UNK token rates are high, the solution might be orthogonal to our approach, requiring further improvement in the base models or tokenizers that could later be integrated with LinguAlchemy.
@@ 5.2 Effect of Scaling URIEL Loss
The classification and URIEL losses operate on different scales. Simply adding these losses together would cause the model to prioritize the loss with the larger magnitude. During the early stages of training, we observe that the classification loss is approximately ten times larger than the URIEL loss. In this section, we explore the impact of various scaling factors applied to the URIEL loss.
Figure 3: Average performance of unseen languages under various URIEL loss scaling factors.
@@
We investigate the effect of consistently scaling the URIEL loss by different factors. The results are illustrated in Figure 3. Notably, because we utilize the scale-invariant optimizer AdamW, there is no risk of gradients becoming excessively large due to high loss values.
As observed in Figure 3, for mBERT, the performance on unseen languages improves significantly when scaling the URIEL loss by 10x, achieving the highest performance of 68.68%. However, as the scaling factor increases further (e.g., 25x, 50x, and 100x), the performance starts to decline, indicating that overly emphasizing the URIEL loss has diminishing returns or even negative effects. Without any scaling (0x), the model performs poorly, showing the importance of scaling the URIEL loss.
For XLM-R, the trend is more stable. The performance fluctuates slightly across different scaling factors but remains generally consistent, with the highest performance achieved at the 10x scaling factor (80.32%). Larger scaling factors (e.g., 50x, 100x) do not lead to substantial improvements and may even cause minor drops in performance. This suggests that while scaling helps, XLM-R is less sensitive to the URIEL loss scaling than mBERT.
Overall, a scaling factor of 10$\times$ appears to give the best balance between classification and URIEL losses.
@@
Introducing a scaling factor adds another tunable hyperparameter, which can complicate the training process. Ideally, we seek a balanced weighting between the classification and URIEL losses. Instead of exhaustively testing various scaling factors, an adaptive scaling approach is more cost-effective and advantageous. Here, we explore two strategies: dynamic scaling and trainable scaling factors. The results of these approaches are presented in Table 4.
@@
Interestingly, these dynamic scaling methods do not significantly outperform a constant scaling factor. Specifically, a 10$\times$ scaling achieves the best performance for mBERT, while dynamic scaling only marginally outperforms the 10$\times$ scaling for XLM-R. Therefore, in scenarios with limited computational resources, a 10$\times$ scaling factor is recommended. However, with more computational capacity, exploring different scaling factors may yield marginal gains.
[…]
@@ 5.4 Seen Language Performance
In 4, we compare the performance of the BERT and XLMR models across different language families. The dotted lines represent the language families that were included in training (i.e., seen languages), and the solid gray lines represent language families that were unseen in training. Notably, variance for unseen languages is indicated by the shaded areas. This figure shows that while both models generally improve across training groups, the performance of certain language families, particularly seen languages, varies. For instance, in the case of the BERT model, languages from the Sino-Tibetan family demonstrate relatively poor performance across training stages compared to other families, even though they are part of the seen group. Meanwhile, XLMR shows a more consistent performance boost across language families, including unseen ones, though Indo-European languages perform better overall.
@@ 6 Conclusion
We introduced LinguAlchemy, a novel approach that demonstrates strong performance across 30+ unseen languages on intent classification and semantic relatedness tasks. Our method hinges on the integration of linguistic knowledge through the URIEL vectors, enhancing the language model’s ability to generalize across a diverse set of languages. We also proposed AlchemyScale and AlchemyTune, which employs a hyperparameter search for the URIEL scaling factor. This is achieved by two key strategies: (1) weight-averaging classification and URIEL loss, and (2) learning to balance the scale between classification and URIEL loss. LinguAlchemy achieves a massive performance improvement on low-resource languages in multiple downstream tasks including intent classification ($\uparrow 18.50$), news classification ($\uparrow 31.46$), and semantic relatedness ($\uparrow 0.32$).
@@ 4 Sino-Tibetan Train Group 3 + my-MM, zh-CN, zh-TW 36
Appendix B Language Family Experiment
@@
Appendix C Appendix: Language Identification (LID) Experiments